It is estimated that 25% of all laptops of a brand are sent to service within two years. Among these, sixty percent are repaired and the other forty percent should be replaced with new ones. If a university buy 10 of these laptops, calculate the probability that exactly 2 laptops are replaced in two years?
Answer #1

Solution

Given that ,

p = 0.25

1 - p = 1 - 0.25 = 0.75

n = 10

Using binomial probability formula ,

P(X = x) = (n C x) * px * (1 - p)n - x

P(X = 2 ) = (10C 2) * (0.25)2 * (0.75)8

= 0.281567

Probability = 0.2816